# GPU-Accelerated Routability-Driven Automated Placement for Large-Scale Photonic Integrated Circuits

### v1

@@ 3.6. Legalization
Since the components in the PICs have a relatively large feature size, we directly perform macro legalization after the global placement. The greedy Macro Legalization operator legalizes movable macros in two stages, first a coarse Hannan-grid pass for small clusters and blocked macros, then an LP/graph-based refinement, alternating until total and weighted displacement are minimized and the best legal layout is retained.
@@ 4. Evaluation Results
The Apollo framework builds upon DREAMPlace, with PIC-specific optimization. All the experiments are conducted on AMD EPYC 7763 Linux server with a 2.9GHz CPU and NVIDIA RTX A6000 GPU.
Figure 6. Proper placement constraints ensure high PIC layout quality on ADEPT_16x16 circuit.
@@
Evaluation Metrics. We evaluate the placement solution quality using several key metrics: routability, number of crossings (#CR), and runtime. Routability is defined as the ratio of successfully routed nets to the total number of nets. To evaluate routability, we perform exact waveguide routing on the legalized placement using an open-source PIC router LiDAR (29). Crossing counts, as an indirect routability indicator, are computed directly from the placement by detecting intersections of the straight-line segments representing each two-pin net. Insertion loss $IL_{max}$ is a critical metric related to waveguide bending angles, routing length, and crossings that implies the optical signal integrity. Note that without a fully DRV-free layout, it is not meaningful to evaluate post-routing insertion loss. Instead, we estimate pre-routing insertion loss $IL_{max}$ after the placement stage using the estimated routing length and port orientations to predict bending angles, then applying a waveguide insertion loss model.
@@ 4.1. Handling of Design Constraints
Constraints are absolutely critical in PIC placement. Due to the complex port alignments in PIC circuits, unreasonable device placement can severely degrade routability, make it impossible to produce a legal layout, induce thermal crosstalk, and exacerbate manufacturing variability. Therefore, it is essential to impose layout constraints from the very beginning. Figure 6 illustrates how placement evolves as left-alignment and y-center-alignment constraints are gradually applied. With no constraints, components are placed in a highly irregular fashion, leading to severe port misalignment and an entirely unrouteable solution. As constraints accumulate, the circuit becomes increasingly regular: grouping certain components for left alignment arranges them into neat columns, though the groups remain too close and can introduce crosstalk; adding y-center alignment then transforms the layout into a true grid, perfectly aligning ports and markedly enhancing routability. To ensure a fair comparison in our experiments, we apply the same constraint set to all baseline methods so that each can produce a valid, reasonable placement.
@@ 4.2. Main Results
Clements_8x8_S 7 43.68% 30.15 0 72.41% 15.80 0 98.85% 21.13
Clements_8x8_L 8 50.42% 31.64 0 77.73% 16.48 0 98.85% 23.07
Clements_16x16_S 20 63.70% 43.67 13 41.91% 22.33 2 95.05% 31.89
Clements_16x16_L 28 59.43% 45.19 17 48.64% 23.01 2 96.04% 32.64
ADEPT_8x8_S 38 48.82% 39.64 61 42.73% 21.58 28 90.05% 22.79
ADEPT_8x8_L 49 70.63% 40.92 53 61.85% 22.63 33 95.04% 23.11
ADEPT_16x16_S 83 47.96% 61.57 105 35.42% 34.36 79 98.12% 35.88
ADEPT_16x16_L 124 40.87% 63.24 113 33.79% 35.62 87 90.60% 39.62
ADEPT_32x32_S 261 44.20% 117.41 209 62.71% 64.43 206 96.74% 65.99
ADEPT_32x32_L 263 48.37% 124.73 253 46.74% 69.89 237 93.38% 71.62
ADEPT_64x64_S 381 43.89% 235.31 358 41.82% 133.39 314 97.10% 105.96
ADEPT_64x64_L 303 48.17% 238.47 297 50.83% 136.71 209 87.62% 110.93
@@
Figure 7. Impact of the exponent $\alpha$ in our cosWA wirelength on routability on ADEPT_16x16.
@@
Clements_16x16_S 2 95.05% 2 95.05%
ADEPT_32x32_S 206 96.74% 183 95.36%
ADEPT_64x64_S 314 97.10% 304 96.32%
Figure 8. Visual comparison of post-routing layout (viewed in KLayout) generated by Cypress and our Apollo.
For the parameters in the proposed cosWA function, we first performed a parameter sweep on the Clements_16x16 benchmark. As shown in Fig. 7, we varied $\alpha$ from 1.0 to 2.0 in increments of 0.1, ran placement followed by full routing at each setting, and recorded the resulting routability. We then selected $\alpha=1.4$, the value that produced the highest routability for all subsequent experiments. As for the margin parameter $c$ in cosWA, because we employ Manhattan routing in later stages, we set c=0; that is, once the ports achieve a 90-degree orientation, no further bending-angle penalty is applied.
Moreover, in all subsequent experiments, we initialize every component at a random center location. As Table  2 shows, even when compared to manually placing components in favorable starting positions, our method still converges to high-quality solutions under random center initialization, demonstrating the robustness of the proposed approach.
Table 1 shows that Apollo consistently outperforms two baselines, four Clements and four ADEPT benchmarks, under both compact relaxed die sizes. Apollo achieves a geometric mean routability of 94.79%, versus 50.85% for DREAMPlace and 51.38% for Cypress. In all S-size runs, Apollo exceeds 98.5% routability, drives net crossings down to zero on Clements_8$\times$8_S and Clements_16$\times$16_S, and finishes in under 25s on average, roughly 1.8$\times$ faster than DREAMPlace and on par with Cypress. Even on the largest ADEPT_64x64_L circuit, Apollo sustains 87.6% routability (vs. $\sim$48%/51%), reduces crossings by $\sim$30%. Unlike the baselines, which optimize a generic VLSI wirelength objective, Apollo is designed specifically for PIC placement and explicitly accounts for the spatial footprint of curved waveguide bends and crossings. This routing-aware formulation yields post-placement layouts that are inherently highly routable. In contrast, DREAMPlace+RO and Cypress, only produce layouts that succeed in only 50.85% and 51.38% of total nets, respectively. Even though DREAMPlace+RO performs node inflation to create extra port spacing based on congestion maps, it does not capture the area overhead of curvy bends, leaving many nets difficult to access and route.
@@ 4.3.1. Wirelength Function Comparison
Clements_8x8_S 1.64E+04 0.32 1.57E+04 0.30 1.66E+04 0.33 1.57E+04 0.31
Clements_16x16_S 5.61E+04 0.59 8.24E+04 0.73 5.63E+04 0.60 5.37E+04 0.60
ADEPT_8x8_S 2.44E+04 0.61 3.09E+04 0.67 2.40E+04 0.60 2.06E+04 0.57
ADEPT_16x16_S 5.96E+04 1.31 7.49E+04 1.37 5.62E+04 1.28 5.17E+04 1.25
ADEPT_32x32_S 1.63E+05 2.58 1.74E+05 2.46 1.60E+05 2.44 1.17E+05 2.32
ADEPT_64x64_S 2.74E+05 3.30 3.17E+05 3.43 2.96E+05 3.04 2.44E+05 2.89
@@
To evaluate the impact of different wirelength models on photonic layout quality, we compare weighted-average wirelength (WA-WL) (10), log-sum-exponential (LSE) (12), quadratic wirelength, and our proposed cosine-weighted wirelength (cosWA). As shown in Table 3, cosWA achieves the lowest maximum insertion loss ($IL_{\max}$), outperforming WA-WL, LSE, and quadratic by 1.10$\times$, 1.13$\times$, and 1.05$\times$, respectively. It also produces the smallest total bending angle improvements of roughly 1.18$\times$, 1.38$\times$, and 1.21$\times$ over WA-WL, LSE, and quadratic wirelength. These reductions in loss and detour translate directly into compact layouts, reduced local congestion, and ultimately significantly improved routability.
@@ 4.3.2. Spacing Model Comparison
Table 4. Routing success rates (higher the better) under different placement spacing models (w/o spacing, port-density-based component inflation($PI$) and our spacing model with different configuration) on various benchmarks (S is small chip space and L is large chip space). Our net spacing model, based on joint waveguide, bending, and crossing modeling (last column), leads to the best routing success rate. PI: cell inflation based on pin density.
Benchmark w/o spacing $PI$ Only basic spacing ($P_{\text{dens}}$) $P_{\text{dens}}+R_{\text{cong}}$
@@
Clements_8x8_S 94.25% 95.40% 95.40% 94.25% 98.85%
Clements_8x8_L 93.01% 96.43% 95.54% 97.70% 98.85%
Clements_16x16_S 82.51% 85.81% 85.48% 87.13% 95.05%
Clements_16x16_L 74.59% 85.15% 87.13% 93.07% 96.04%
ADEPT_8x8_S 76.38% 78.74% 70.87% 87.40% 90.05%
ADEPT_8x8_L 65.80% 70.08% 76.38% 87.74% 95.04%
ADEPT_16x16_S 60.50% 61.76% 54.23% 95.61% 98.12%
ADEPT_16x16_L 63.95% 69.91% 84.01% 86.21% 90.60%
ADEPT_32x32_S 49.15% 51.89% 53.85% 96.35% 96.74%
ADEPT_32x32_L 62.40% 63.31% 72.62% 83.96% 93.38%
ADEPT_64x64_S 40.87% 52.65% 66.44% 91.29% 97.10%
ADEPT_64x64_L 50.98% 66.32% 70.77% 76.94% 87.62%
@@
Clements_8x8_S Success 1.56E+04 0.294 Success 1.59E+04 0.357 Diverge 1.61E+04 0.307 Diverge 1.53E+04 0.285
Clements_16x16_S Success 5.75E+04 0.684 Success 5.35E+04 0.655 Diverge 5.54E+04 0.666 Success 5.96E+04 0.690
ADEPT_8x8_S Success 1.78E+04 0.464 Success 1.81E+04 0.467 Success 1.95E+04 0.724 Diverge NA NA
ADEPT_16x16_S Success 4.18E+04 1.097 Success 4.59E+04 1.610 Success 4.62E+03 1.540 Success 4.42E+04 1.363
ADEPT_32x32_S Success 1.03E+05 1.497 Success 1.20E+05 2.958 Diverge 1.15E+05 1.932 Diverge NA NA
ADEPT_64x64_S Success 2.48E+05 1.901 Success 2.52E+05 2.619 Success 2.69E+05 3.779 Success 2.93E+05 2.706
@@
Although cosWA wirelength substantially reduces unnecessary waveguide detours, regions traversed by multiple waveguides can still become heavily congested. This is especially true for multi-port components, where servicing or escaping several ports in a tight area entails numerous bends and crossings—scenarios that conventional VLSI node-inflation cannot adequately resolve. To address this, we borrow from automated waveguide routing practices and introduce explicit spacing models to mitigate congestion arising from both bends and crossings. We conduct an ablation study comparing: (1) no spacing mode, (2) port-density inflation ($PI$): adding Halo to cells proportional to the product of number of ports and bending radius for the cell, our proposed basic net spacing schemes, including two variants: (3) bend-radius spacing ($r_{\text{bend}}$) and (4) the port-count-crossing product $S_{\text{crs}}$, and (5) our final net spacing model with basic spacing and congestion cores ($P_{\text{dens}}+R_{\text{cong}}$).
As Table 4 shows, port-density-based cell inflation does improve routability over no spacing model (geometric-mean success rate rises from 67.87% to 73.12%), but it still fails on ADEPT circuits with heavy multi-port crossings. Introducing a basic bend-radius clearance ($r_{\rm bend}$) further boosts the geo-mean to 76.06%, yet it still overlooks congestion at port crossings. When we explicitly account for crossing-induced spacing, the routability jumps to 89.80%. Finally, our full net spacing model, combining port-density and crossing-aware spacing, delivers an average success rate of 94.79% across all benchmarks.
@@ 4.3.3. Optimizer Comparison
Table 5 compares our proposed BNAG optimizer with NAG+BB (3), NAG (16), and Adam (13) over six benchmarks under the same number of iterations. While NAG and Adam optimizers diverge on several test cases, both BNAG and NAG+BB converge across all benchmarks. Moreover, BNAG achieves an improvement of approximately 1.04×1.04$\times$1.04 × in total bending cost and 1.29×1.29$\times$1.29 × in maximum insertion loss.

### v2

@@ 3.6. Legalization
Since the components in the PICs have a relatively large feature size, we directly perform macro legalization (16) after the global placement. The greedy Macro Legalization operator legalizes movable macros in two stages: first, a coarse Hanan-grid pass for small clusters and blocked macros, and then an LP/graph-based refinement, alternating until total and weighted displacement are minimized and the best legal layout is retained.
@@ 4. Evaluation Results
The Apollo framework builds upon DREAMPlace, with PIC-specific optimization. All the experiments are conducted on an AMD EPYC 7763 Linux server with a 2.9GHz CPU and NVIDIA RTX A6000 GPU.
@@
Benchmarks. In our customized realistic benchmark suites, we use the Clements-style Mach-Zehnder Interferometer (MZI) (18) arrays and an auto-searched photonic tensor core design (ADEPT) (8) at two different scales. For each circuit, we generate two chip configurations: $S$ size with a very compact die size with a 5 $\mu m$ bending radius and $L$ size with a more relaxed die size with a 10 $\mu m$ bending radius, and use a $\mu m ×\times10$$\mu m$ crossing size.
@@
Evaluation Metrics. We evaluate the placement solution quality using several key metrics: routability, number of crossings (#CR), and runtime. Routability is defined as the ratio of successfully routed nets to the total number of nets. To evaluate routability, we perform exact waveguide routing on the legalized placement using an open‑source PIC router LiDAR (29). Crossing counts, as an indirect routability indicator, are computed directly from the placement by detecting intersections of the straight‑line segments representing each two‑pin net. Insertion loss $IL_{max}$ is a critical metric related to waveguide bending angles, routing length, and crossings that implies the optical signal integrity. Note that without a fully DRV-free layout, it is not meaningful to evaluate post-routing insertion loss. Instead, we estimate the pre-routing insertion loss $IL_{max}$ after the placement stage from the estimated routing length and port orientations, predicting bending angles and applying an analytical waveguide insertion loss model (29).
@@ 4.1. Handling of Design Constraints
Constraints are absolutely critical in PIC placement. Due to the complex port alignments in PIC circuits, unreasonable device placement can severely degrade routability, make it impossible to produce a legal layout, induce thermal crosstalk, and exacerbate manufacturing variability. Therefore, it is essential to impose layout constraints from the very beginning. Figure 6 illustrates how placement evolves as left‑alignment and y‑center‑alignment constraints are gradually applied. With no constraints, components are placed in a highly irregular fashion, leading to severe port misalignment and an entirely unrouteable solution. As constraints accumulate, the circuit becomes increasingly regular: grouping certain components for left alignment arranges them into neat columns, though the groups remain too close and can introduce crosstalk; adding y‑center alignment then transforms the layout into a true grid, perfectly aligning ports and markedly enhancing routability. To ensure a fair comparison in our experiments, we apply the same constraint set to all baseline methods so that each can produce a valid, reasonable placement.
@@ 4.2. Main Results
[…]
@@
[…]
@@
[…]
For the parameters in the proposed cosWA function, we first performed a parameter sweep on the ADEPT_16x16 benchmark. As shown in Fig. 7, we varied $\alpha$ from 1.0 to 2.0 in increments of 0.1, ran placement followed by full routing at each setting, and recorded the resulting routability. We then selected $\alpha=1.4$, the value that produced the highest routability for all subsequent experiments. As for the margin parameter $c$ in cosWA, because we employ Manhattan routing in later stages, we set c=0; that is, once the ports achieve a 90-degree orientation, no further bending‐angle penalty is applied.
Moreover, in all subsequent experiments, we initialize every component at a random center location. As Table 2 shows, even when compared to manually placing components in favorable starting positions, our method still converges to high‑quality solutions under random center initialization, demonstrating the robustness of the proposed approach.
[…]
Unlike the baselines, which optimize a generic VLSI wirelength objective, Apollo is designed specifically for PIC placement and explicitly accounts for the spatial footprint of curved waveguide bends and crossings. This routing‑aware formulation yields post‑placement layouts that are inherently highly routable. Even though DREAMPlace+RO performs node inflation to create extra port spacing based on congestion maps, it does not capture the area overhead of curvy bends, leaving many nets difficult to access and route.
@@ 4.3.1. Wirelength Function Comparison
[…]
@@
[…]
@@ 4.3.2. Spacing Model Comparison
Table 4. Routing success rates (higher the better) under different placement spacing models (w/o spacing, port-density-based component inflation (PIPI) and our spacing model with different configuration) on various benchmarks (S is small chip space and L is large chip space). Our net spacing model, based on joint waveguide, bending, and crossing modeling (last column), leads to the best routing success rate. PI: cell inflation based on pin density.
Benchmark w/o spacing PIPI Only basic spacing ($P_{\text{dens}}$) $P_{\text{dens}}+R_{\text{cong}}$
@@
[…]
@@
[…]
@@
Although cosWA wirelength substantially reduces unnecessary waveguide detours, regions traversed by multiple waveguides can still become heavily congested. This is especially true for multi‑port components, where servicing or escaping several ports in a tight area entails numerous bends and crossings—scenarios that conventional VLSI node‑inflation cannot adequately resolve. To address this, we borrow from automated waveguide routing practices and introduce explicit spacing models to mitigate congestion arising from both bends and crossings. We conduct an ablation study comparing: (1) no spacing mode, (2) port-density inflation ($PI$): adding Halo to cells proportional to the product of number of ports and bending radius for the cell, our proposed basic net spacing schemes, including two variants: (3) bend-radius spacing ($r_{\text{bend}}$) and (4) the port-count-crossing product $S_{\text{crs}}$, and (5) our final net spacing model with basic spacing and congestion cores ($P_{\text{dens}}+R_{\text{cong}}$).
As Table 4 shows, port‑density-based cell inflation does improve routability over no spacing model (geometric‑mean success rate rises from 67.87% to 73.12%), but it still fails on ADEPT circuits with heavy multi‑port crossings. Introducing a basic bend‑radius clearance ($r_{\rm bend}$) further boosts the geo‑mean to 76.06%, yet it still overlooks congestion at port crossings. When we explicitly account for crossing‑induced spacing, the routability jumps to 89.80%. Finally, our full net spacing model, combining port‑density and crossing-aware spacing, delivers an average success rate of 94.79% across all benchmarks.
@@ 4.3.3. Optimizer Comparison
Table 5 compares our proposed BNAG optimizer with NAG+BB (3), NAG (16), and Adam (13) over six benchmarks under the same number of iterations. While NAG and Adam optimizers diverge on several test cases, both BNAG and NAG+BB converge across all benchmarks. Moreover, BNAG achieves an improvement of approximately $1.04\times$ in total bending cost and $1.29\times$ in maximum insertion loss.